# Few-Shot Video Object Segmentation in X-Ray Angiography Using Local Matching and Spatio-Temporal Consistency Loss

### v1

@@ Abstract
High-quality, densely annotated data serve as a crucial foundation for developing robust X-ray angiography segmentation models. However, obtaining per-object pixel-level annotations in the medical domain is both expensive and time-consuming, often requiring close collaboration between clinical experts and developers. This paper aims to reduce the annotation costs of X-ray angiography videos by leveraging few-shot video object segmentation (FSVOS), which separates target objects from the background using only a single annotated frame during inference. We introduce a novel FSVOS model that employs a local matching strategy to restrict the search space to the most relevant neighboring pixels. Rather than relying on inefficient standard im2col-like implementations (e.g., spatial convolutions, depthwise convolutions and feature-shifting mechanisms) or hardware-specific CUDA kernels (e.g., deformable and neighborhood attention), which often suffer from limited portability across non-CUDA devices, we reorganize the local sampling process through a direction-based sampling perspective. Specifically, we implement a non-parametric sampling mechanism that enables dynamically varying sampling regions. This approach provides the flexibility to adapt to diverse spatial structures without the computational costs of parametric layers and the need for model retraining. To further enhance feature coherence across frames, we design a supervised spatio-temporal contrastive learning scheme that enforces consistency in feature representations. In addition, we introduce a publicly available benchmark dataset for multi-object segmentation in X-ray angiography videos (MOSXAV), featuring detailed, manually labeled segmentation ground truth. Extensive experiments on the CADICA, XACV, and MOSXAV datasets show that our proposed FSVOS method outperforms current state-of-the-art video segmentation methods in terms of segmentation accuracy and generalization capability (i.e., seen and unseen categories). This work offers enhanced flexibility and potential for a wide range of clinical applications.
@@ 1 Introduction
Driven by the paradigm of class-agnostic mask tracking, the computer vision community has increasingly focused on few-shot video object segmentation (FSVOS) [1, 2, 3, 4, 5, 6, 7]. This shift prioritizes robust tracking mechanisms over the explicit learning of object-specific semantic representations, thereby mitigating the heavy reliance on large-scale annotated datasets. The goal of FSVOS is to segment target objects throughout an entire video sequence based on an initial object mask, which can be either manually provided or automatically generated from any single frame. As a widely accepted solution for FSVOS, matching-based methods [3, 4, 5] perform a matching strategy to generate dense correspondences between the query frame and the support frame(s), where the past frames as well as corresponding masks are stored in an external memory to build a feature memory [8]. In Space-Time Memory (STM) network [3], cross-image correspondences is the key component that makes it superior in performance and simplicity. In the pixel-wise retrieval phase, it enables the global content-dependent interactions among different image regions to model long-range dependencies. Through the visualization of dense space-time correspondence, we observe the crucial role of local contexts, such as the neighborhood pixels [9, 10, 11, 12, 13], in establishing correspondences for cross-frame matching (see the similarity scores in Fig. 1). Meanwhile, global and fine-grained dependencies become less critical due to spatial redundancy and noise as the number of features involved in non-local matching increases, particularly in cases of locally recurring motion such as cardiac and respiratory movements, or within specific regions of interest in X-ray angiography videos. On the other hand, matching-based methods [3, 4, 5, 14] focus only on discovering space-time correspondences from inter-frames, while ignoring the valuable temporal consistency of object representation across multiple frames. Indeed, due to only mining the matched pixels across the memory frames, some background pixels/patches are wrongly recognized as highly correlated to the query primary objects. Therefore, it is necessary to make full use of inter-frame consistency to make up for the drawbacks caused by ignoring temporal coherence.
@@ 2.1 Vessel and catheter segmentation
In addition, automatic catheter segmentation has emerged as another active topic in medical image analysis. Existing work has reported catheter segmentation in X-ray images, including electrophysiology (EP) electrodes [24] and EP catheters [25, 26]. More recently, Kim et al. [27] addressed the challenge of detecting active acoustic catheters (AACs) in echocardiography. Their method first applies U-Net to segment the left ventricle, then filters the color response generated by the AAC, and finally performs thresholding to identify the catheter. Yang et al. [28] proposed a 3D ultrasound catheter segmentation method based on Shared-ConvNet, which employs CNNs with shared weights across imaging planes and performs voxel-wise binary classification. Overall, beyond our prior work, the literature on catheter segmentation in AI-driven, ultrasound-guided minimally invasive endovascular surgery (MIES) remains limited. Nguyen et al. [29] introduced an end-to-end, real-time deep learning framework for endovascular intervention, incorporating a novel flow-guided warping function to enforce frame-to-frame temporal continuity. To capture such sequential behavior, Ranne et al. [30] proposed the Attention-in-Attention ResNet for Segmentation (AiAReSeg), which aggregates information across image sequences to infer knowledge about the current frame.
@@ 2.2 Few-shot video object segmentation
FSVOS [1] involves limited human inspection (typically provided in the first frame) to input the segmentation mask for the objects of interest. The methods then segment the desired objects in the remaining frames. Here, few-shot refers to the level of human interaction at test time, not during the training phase. The few-shot methods still rely on the supervised learning paradigm to train a pixel-wise tracking [3, 4, 5, 14] or mask propagation framework. Early few-shot methods [1, 31] employ online fine-tuning on the basis of this supervision, which suffered from high test runtime and were gradually phased out. The follow-up researches have been explored including embedding learning [32, 33], propagation-based [2, 34, 35] and tracking-based [36, 37]. These methods perform frame-to-frame propagation or global matching with the first reference frame, while the context is still limited and it becomes harder to match as the video progresses. To fully exploit the spatial-temporal context over longer distances, recent state-of-the-art methods define the past frames with object masks as feature memory and the current frame as the query [3, 38, 39]. As a representative work, STM [3] utilizes an external memory to store the past frames as well as corresponding masks, where a dense matching is adopted to establish long-range dependencies in order to achieve pixel classification and objects segmentation. Staring from STM [3], matching-based paradigm was adopted and has been extended by many follow-up works [40, 41, 4, 42, 43, 5, 14], which achieve leading accuracy on most benchmarks duo to long-range context support.
@@ 4 Experiments
4.1 Experimental setup
4.1.1 Datasets
4.1.2 Evaluation metric
4.1.3 Implementation details
[…]
4.2 Main Results
@@ 4.3 Ablation Study
4.3.1 Training objective
4.3.2 Sampling strategies
Compared with the baseline, our local sampling module implemented via the unfold sampling function (Eq. 7) improves segmentation performance by 1.8% in terms of mean $\mathcal{J}$. More importantly, it increases inference speed to 25 FPS on the GPU (vs. 20 for the baseline) and to 0.301 FPS and 0.293 FPS on the Intel® Core™ i9-13900H CPU and the AMD Ryzen™ 9 7940HS CPU, respectively (vs. 0.289 and 0.283 for the baseline). To improve efficiency, we first partition the image into patches, similar to the patch embedding module in ViT [21], and then compute similarity within each local window. The main motivation for this design is that applying im2col-like local sampling individually to every pixel is prohibitively expensive. Consistent with this observation, the feature-shift operation based on an im2col-like function not only yields lower segmentation accuracy but also incurs higher computational cost. For fixed and deformable convolution-based sampling, the learnable bias terms fail to capture effective inductive biases due to the depth-wise nature of these operations. In our implementation, we fix the kernel weights to 1 and optimize only the bias parameters. Compared with the baseline, these two sampling operations improve accuracy by 0.5% and 1.0%, respectively, but both result in lower inference speed on both GPU and CPU. In contrast, the 2D neighborhood attention operation achieves strong performance in terms of both accuracy and computational efficiency; however, its deployment, similar to deformable convolution, relies on device-specific CUDA implementations.
@@ 5 Conclusion
Despite these methodological advances, translating a retrospective research prototype into a real-world clinical application remains challenging, particularly with respect to data privacy, system compatibility, and regulatory approval. To bridge this gap, our future work will focus on two fronts: (1) Clinical Prototyping: We aim to develop a DICOM-compatible plugin via open-source platforms like 3D Slicer, enabling intuitive clinician interaction. (2) Regulatory Approval: We will further improve inference latency and data security protocols to meet regulatory standards. We believe this work establishes a robust foundation for “human-in-the-loop” AI systems within standard diagnostic workflows.

### v2

# Lin Xi1,  Yingliang Ma1,2,,  Xiahai Zhuang3

@@ Abstract
High-quality, densely annotated data serve as a crucial foundation for developing robust X-ray angiography segmentation models. However, obtaining per-object pixel-level annotations in the medical domain is both expensive and time-consuming, often requiring close collaboration between clinical experts and developers. This paper aims to reduce the annotation costs of X-ray angiography videos by leveraging few-shot video object segmentation (FSVOS), which separates target objects from the background using only a single annotated frame during inference. We introduce a novel FSVOS model that employs a local matching strategy to restrict the search space to the most relevant neighboring pixels. Rather than relying on inefficient standard im2col-like implementations (e.g., spatial convolutions, depthwise convolutions and feature-shifting mechanisms) or hardware-specific CUDA kernels (e.g., deformable and neighborhood attention), which often suffer from limited portability across non-CUDA devices, we reorganize the local sampling process through a direction-based sampling perspective. Specifically, we implement a non-parametric sampling mechanism that enables dynamically varying sampling regions. This approach provides the flexibility to adapt to diverse spatial structures without the computational costs of parametric layers and the need for model retraining. To further enhance feature coherence across frames, we design a supervised spatio-temporal contrastive learning scheme that enforces consistency in feature representations. In addition, we introduce a publicly available benchmark dataset for multi-object segmentation in X-ray angiography videos (MOSXAV), featuring detailed, manually labeled segmentation ground truth. Extensive experiments on the CADICA, XACV, and MOSXAV datasets show that our proposed FSVOS method outperforms current state-of-the-art video segmentation methods in terms of segmentation accuracy and generalization capability (i.e., seen and unseen categories). This work offers enhanced flexibility and potential for a wide range of clinical applications. Code is available at: https://github.com/xilin-x/XRAVOS
@@ 1 Introduction
Driven by the paradigm of class-agnostic mask tracking, the computer vision community has increasingly focused on few-shot video object segmentation (FSVOS) [1, 2, 3, 4, 5, 6, 7]. This shift prioritizes robust tracking mechanisms over the explicit learning of object-specific semantic representations, thereby mitigating the heavy reliance on large-scale annotated datasets. The goal of FSVOS is to segment target objects throughout an entire video sequence based on an initial object mask, which can be either manually provided or automatically generated from any single frame. As a widely accepted solution for FSVOS, matching-based methods [3, 4, 5] perform a matching strategy to generate dense correspondences between the query frame and the support frame(s), where the past frames as well as corresponding masks are stored in an external memory to build a feature memory [8]. In Space-Time Memory (STM) network [3], cross-image correspondences is the key component that makes it superior in performance and simplicity. In the pixel-wise retrieval phase, it enables the global content-dependent interactions among different image regions to model long-range dependencies. Through the visualization of dense space-time correspondence, we observe the crucial role of local contexts, such as the neighborhood pixels [9, 10, 11, 12, 13], in establishing correspondences for cross-frame matching (see the similarity scores in Fig. 1). Meanwhile, compared to local-side computing solutions [14], global and fine-grained dependencies become less critical due to spatial redundancy and noise as the number of features involved in non-local matching increases, particularly in cases of locally recurring motion such as cardiac and respiratory movements, or within specific regions of interest in X-ray angiography videos. On the other hand, matching-based methods [3, 4, 5, 15] focus only on discovering space-time correspondences from inter-frames, while ignoring the valuable temporal consistency of object representation across multiple frames. Indeed, due to only mining the matched pixels across the memory frames, some background pixels/patches are wrongly recognized as highly correlated to the query primary objects. Therefore, it is necessary to make full use of inter-frame consistency to make up for the drawbacks caused by ignoring temporal coherence.
@@ 2.1 Vessel and catheter segmentation
In addition, automatic catheter segmentation has emerged as another active topic in medical image analysis. Existing work has reported catheter segmentation in X-ray images, including electrophysiology (EP) electrodes [25] and EP catheters [26, 27, 28]. More recently, Kim et al. [29] addressed the challenge of detecting active acoustic catheters (AACs) in echocardiography. Their method first applies U-Net to segment the left ventricle, then filters the color response generated by the AAC, and finally performs thresholding to identify the catheter. Yang et al. [30] proposed a 3D ultrasound catheter segmentation method based on Shared-ConvNet, which employs CNNs with shared weights across imaging planes and performs voxel-wise binary classification. Overall, beyond our prior work, the literature on catheter segmentation in AI-driven, ultrasound-guided minimally invasive endovascular surgery (MIES) remains limited. Nguyen et al. [31] introduced an end-to-end, real-time deep learning framework for endovascular intervention, incorporating a novel flow-guided warping function to enforce frame-to-frame temporal continuity. To capture such sequential behavior, Ranne et al. [32] proposed the Attention-in-Attention ResNet for Segmentation (AiAReSeg), which aggregates information across image sequences to infer knowledge about the current frame.
@@ 2.2 Few-shot video object segmentation
FSVOS [1] differs from fully automated video segmentation [33, 34, 35] by incorporating human input, specifically a first-frame segmentation mask of the target objects. The methods then segment the desired objects in the remaining frames. Here, few-shot refers to the level of human interaction at test time, not during the training phase. The few-shot methods still rely on the supervised learning paradigm to train a pixel-wise tracking [3, 4, 5, 15] or mask propagation framework. Early few-shot methods [1, 36] employ online fine-tuning on the basis of this supervision, which suffered from high test runtime and were gradually phased out. The follow-up researches have been explored including embedding learning [37, 38], propagation-based [2, 39] and tracking-based [40, 41]. These methods perform frame-to-frame propagation or global matching with the first reference frame, while the context is still limited and it becomes harder to match as the video progresses. To fully exploit the spatial-temporal context over longer distances, recent state-of-the-art methods define the past frames with object masks as feature memory and the current frame as the query [3, 42]. As a representative work, STM [3] utilizes an external memory to store the past frames as well as corresponding masks, where a dense matching is adopted to establish long-range dependencies in order to achieve pixel classification and objects segmentation. Staring from STM [3], matching-based paradigm was adopted and has been extended by many follow-up works [43, 4, 5, 15], which achieve leading accuracy on most benchmarks duo to long-range context support.
@@ 3.2.3 Local sampling
In summary, standard im2col-like implementations (e.g., spatial convolutions, depthwise convolutions and feature-shifting mechanisms) are data-rearrangement techniques that transform local neighborhoods into explicit, redundant patches to enable dense matrix multiplication. This process significantly increases memory consumption, as the footprint is proportional to the kernel or window size. In contrast, our direction-based unfolding treats local sampling as a logical offset mapping. By directly computing correlations between query features and sampled feature maps (or via index-based sampling), we eliminate the requirement for an intermediate “flattened” buffer. This results in a more memory-efficient and computationally streamlined implementation that preserves the inductive bias of local sampling without the overhead of physical data duplication.
@@ 4 Experiments
4.1 Experimental setup
4.1.1 Datasets
4.1.2 Evaluation metric
4.1.3 Implementation details
[…]
4.2 Main Results
@@ 4.3.1 Training objective
4.3.2 Sampling strategies
Compared with the baseline, our local sampling module implemented via the unfold sampling function (Eq. 7) improves segmentation performance by 1.8% in terms of mean $\mathcal{J}$. More importantly, it increases inference speed to 25 FPS on the GPU (vs. 20 FPS for the baseline) and to 0.301 FPS and 0.293 FPS on the Intel® Core™ i9-13900H CPU and the AMD Ryzen™ 9 7940HS CPU, respectively (vs. 0.289 and 0.283 for the baseline). To improve efficiency, we first partition the image into patches, similar to the patch embedding module in ViT [22], and then compute similarity within each local window. The main motivation for this design is that applying im2col-like local sampling individually to every pixel is prohibitively expensive. Regarding clinical utility, our current PyTorch implementation requires approximately 3.32 and 3.41 seconds per image on the aforementioned Intel and AMD CPUs. While this execution speed is not yet optimized for real-time intraoperative use, it is well-suited for offline analysis and automated batch processing within the research phase. Our method achieves a performance gain over the global matching baseline while maintaining competitive accuracy, suggesting it can be effectively integrated into clinical workflows that prioritize reliability over latency, such as retrospective cohort studies or pre-procedural planning. Consistent with this observation, we found that the feature-shift operation based on im2col-like functions yield both lower segmentation accuracy and higher computational costs. For fixed and deformable convolution-based sampling, the learnable bias terms fail to capture effective inductive biases due to the depth-wise nature of these operations. In our implementation, we fix the kernel weights to 1 and optimize only the bias parameters. Compared with the baseline, these two sampling operations improve accuracy by 0.5% and 1.0%, respectively, both resulted in reduced inference speeds across GPU and CPU architectures. In contrast, while 2D neighborhood attention achieves strong performance and efficiency, its deployment, much like deformable convolution, remains constrained by a dependency on device-specific CUDA implementations.
@@ 5 Conclusion
Despite these methodological advances, translating a retrospective research prototype into a real-world clinical application presents significant challenges, particularly regarding data privacy, system interoperability, regulatory compliance, and real-time deployment. To bridge this gap, our future work will focus on three strategic fronts: (1) Clinical prototyping: We aim to develop a DICOM-compatible plugin via open-source platforms such as 3D Slicer, facilitating intuitive clinician interaction and seamless data integration. (2) Regulatory and security standards: We will further optimize inference latency and strengthen data security protocols to align with stringent medical device regulatory standards. (3) High-performance deployment: To achieve real-time capability, we will transition from our current Python-based implementation to high-efficiency C++ frameworks, specifically leveraging LibTorch and ONNX Runtime to enable advanced parallel processing and hardware acceleration. We believe this work establishes a robust foundation for “human-in-the-loop” AI systems, ultimately enhancing standard diagnostic workflows with reliable, high-performance assistance.